OpenStudy (anonymous):

2x+4y=28 x-2y=10 using addition

OpenStudy (anonymous):

x = 12 y = 1

OpenStudy (anonymous):

time (2) x-2y=10 you get 2x-4y=20 2x-4y=20 2x+4y=28 Add --------- 4x= 48 x=48/4 x=12 Substitution x=12 find y 2x+4y=28 2(12)+4y=28 24 + 4y = 28 4y= 28-24 y = 4/4 y=1